OpenStudy (anonymous):

sa solution is made by dissolving 26.42 g of (NH4)2SO4 in enough H2O to make 50.00 mL of solution. what is the molarity of the solution

OpenStudy (anonymous):

first find the molar mass of (NH4) 2S04. so. (14+4*1.01)+(2*32.06+8*16) = 210.16 We know know that if we had 210.16 g of NH4 2SO4 we would have a mole. so we do 26.42/210.16 = .1257 moles of NH4 2SO4. 50ml is .05L molarity is moles per liter. so .1257 mol/.05 L = 2.514 Molarity Or M of NH4 2SO4. Listen though. This is wrong if you meant (NH4)subscript2 then SO4. But this should be right if you meant NH4 2SO4 which means there is 2 S and 8 H.

OpenStudy (accessdenied):

I suspect this is (NH4)subscript2 SO4 on two reasons: (1) You don't typically dissolve already-dissociated ions to my knowledge, and (2) The charges conveniently balance out: (+1)*2 + -2 = 0

OpenStudy (accessdenied):

BUT the same process works either way, finding the molar mass of the substance to obtain moles solute. Divide by volume of solvent, to obtain molarity.